OpenStudy (anonymous):

Can someone help me on this In a canoe race, a team paddles downstream 480 m in 60s. The same team makes the trip upstream in 80 s. Find the team's rate in still water and the rate of the current.

OpenStudy (anonymous):

Can you help?

OpenStudy (anonymous):

Still hard to understand...

hero (hero):

Luis, try putting it all on one pic. That would probably help

OpenStudy (anonymous):

Then what is the current speed of the water?

OpenStudy (anonymous):

That kind of makes sense... I think...

OpenStudy (anonymous):

Uh... ok...

hero (hero):

480 = (r + w)60 480 = (r - w)80 480 = 60r + 60w 480 = 80r - 80w w = (480 - 60r)/60 w = (480 - 80r)/-80 w = w -80(480 - 60r) = 60(480 - 80r) -8(48 - 6r) = 6(48 - 8r) -384 + 48r = 288 - 48r 384 + 288 = 96r 672 = 96r 672/96 = r r = 7

hero (hero):

still have to find w

OpenStudy (anonymous):

Thats what I got too!

OpenStudy (anonymous):

Ok?

hero (hero):

The rate is 7 meters per second

OpenStudy (anonymous):

Then wht is the currents speed?!?

hero (hero):

I'll find w

hero (hero):

w = 1 meter per second

OpenStudy (anonymous):

How did you get 1?

hero (hero):

I solved for w

OpenStudy (anonymous):

Well, can you show the steps of how you got it ?

hero (hero):

I plugged it back into one of the original equations I created earlier

hero (hero):

Yup, I know Luis, but you weren't supposed to

OpenStudy (anonymous):

Well thats what I did too, but I kept getting 473...

hero (hero):

473 is definitely not it

OpenStudy (anonymous):

Yah, thats what got me confused...

hero (hero):

d = rt 480 = (r + w)60 480 = (r - w)80 480 = 60r + 60w 480 = 80r - 80w w = (480 - 60r)/60 w = (480 - 80r)/-80 w = w -80(480 - 60r) = 60(480 - 80r) -8(48 - 6r) = 6(48 - 8r) -384 + 48r = 288 - 48r 384 + 288 = 96r 672 = 96r 672/96 = r r = 7 w = 1

hero (hero):

480 = (r + w)60 480 = (7+1)60 480 = (8)60 480 = 480

OpenStudy (anonymous):

Ok, now I just got confused again.

hero (hero):

or you can do this: 480 = (r + w)60 480 = (7 + w)60 480 = 420 + 60w 480 - 420 = 60w 60 = 60w 60/60 = w 1 = w

hero (hero):

If you're still confused, I can't help you.

OpenStudy (anonymous):

Ok now I know my mistake, I forgot to put in the 60...hehe

OpenStudy (anonymous):

Thanks!

hero (hero):

Good luck

OpenStudy (anonymous):

I will need it for sure!